OpenStudy (anonymous):

Evaluate |-4| - |10|

OpenStudy (anonymous):

u think ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I have 2 know what u think so i can help

OpenStudy (anonymous):

it is plz refer to the coc it says not to give direct answers

pooja195 (pooja195):

@montecristo73 here on OS we dont give direct answer :)

OpenStudy (anonymous):

the two lines are called absolute line they make every thing inside them positive so it would be 4-10 which would be -6

OpenStudy (anonymous):

@montecristo73 thats kind of rude it might not be so easy for @jilliantaylorhughes

OpenStudy (anonymous):

My bad. I apologize.

OpenStudy (michele_laino):

please keep in mind that: \[\begin{gathered} \left| { - 4} \right| = 4 \hfill \\ \left| {10} \right| = 10 \hfill \\ \end{gathered} \]

TheSmartOne (thesmartone):

Also note that any number inside of a\(\bf |~|\) will always come out as positive even if the number inside it was negative. For example \(\bf |-4| =4\) \(\bf |-9| =9\) \(\bf |5| =5\) \(\bf |7| =7\)